OpenStudy (dls):

A line passing through (2,2) is perpendicular to the line 3x+2y=3.its y intercept is?

OpenStudy (anonymous):

Since the Lines are perpendicular m 1 * m2 = -1

OpenStudy (dls):

m=2/3

OpenStudy (dls):

kthx

OpenStudy (anonymous):

3x+2y=3 3x + 2y - 3 =0 m1 = -a/b = -3/2 m2 = 2/3 Nw use Point Slope Form

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

Hope nw u can Solve it

OpenStudy (dls):

3y-2x-2=0

OpenStudy (dls):

i got till here before too..im gettig 2/3..itswrrong

OpenStudy (anonymous):

(2,2) y-y1 = m(x-x1) y-2 = 2/3 ( x-2) 3y - 6 = 2x - 4 2x - 3y +2=0

OpenStudy (anonymous):

Nw to Find y intercept.....Put x = 0

OpenStudy (dls):

IM GETTING 2/3 ITS WRONG:/

OpenStudy (anonymous):

So -3y = -2 y = 2/3

OpenStudy (dls):

WRONG

OpenStudy (anonymous):

But that Shuld be it.....Can u Check ur Question and Answer Key Once Agian ?

OpenStudy (dls):

4/3..

OpenStudy (anonymous):

hm..hm..

OpenStudy (anonymous):

Are you 100% sure you copied the correct equation and point? If not, then @Yahoo! is spot on.

OpenStudy (dls):

yes..i just want to confirm is there any hope of getting 4/3 or is there defintely a misprint in my ans key..

OpenStudy (anonymous):

Read the second part of my comment. It already states that if it is the correct equation and point, then it's a typo/misprint.